OpenStudy (anonymous):

Please help. (2^3)^-8 Answer using exponential notation with positive exponents.

OpenStudy (anonymous):

1 / 8^8

OpenStudy (anonymous):

Can you explain how you reached your answer please?

OpenStudy (anonymous):

2^3 = 2*2*2 = 8

OpenStudy (anonymous):

then , 8^-8 means 1 / 8^8

OpenStudy (anonymous):

as an example , 1^-3 would mean 1 / 1^3

OpenStudy (anonymous):

Remember this as a rule

OpenStudy (anonymous):

Wouldn't you multiply the number by itself the number of times of the exponent to clear it?

OpenStudy (anonymous):

Do you want to remove the exponent ?

OpenStudy (anonymous):

If you want to remove the exponent then yes, that is what you would do

OpenStudy (anonymous):

I don't think so, but this stuff confuses me so much that I just don't know. The question in my text is written just as I wrote it here, so I believe it's mainly just asking for simplification, not solving. So, I'm guessing that, no, I don't need to clear the exponents. What is the rule I should remember?

OpenStudy (anonymous):

The rule that you should remember in such a question where you are required to give your answer in positive exponent is that , if for exampe you are given 1^-3 , that equals 1 / 1^3 .It's as simple as that

OpenStudy (anonymous):

When you divide it by 1 , the sign of your exponent changes. When you divided 1^-3 by 1 , the sign of your exponent changed , hence your answer would be 1 / 1^3

OpenStudy (anonymous):

Ok. :Thank you so much. I think I understand a little better now. :)

OpenStudy (anonymous):

np :)